# MunchSonic: Tracking Fine-grained Dietary Actions through Active Acoustic Sensing on Eyeglasses

### v1

@@ 5.1. Evaluation of Frame-Level Inference
We summarize the results of the leave-one-participant-out evaluation of sliding window prediction in Figure 3. According to the evaluation, the average macro F1-score across all 12 participants is 0.935 with a standard deviation of 0.031. It is evident that the MunchSonic system can track fine-grained dietary events with precision and recall of more than 90% for most participants. P08 demonstrates the worst performance in terms of both precision and recall. Additionally, P04 and P13 yielded slightly lower recall compared to the average. Our analysis of the ground truth videos for those participants suggests that the slightly worse performance can be attributed to the participants touching their faces several times during the study for activities not related to eating. This phenomenon is also evident in the confusion matrix in Figure 4. The performance degradation of MunchSonic for face-touch actions can be related to the imbalanced nature of the dataset collected in a completely uncontrained manner, which contains very few samples of this particular activity. Nonetheless, MunchSonic exhibits robust performance across all activities, with a mean precision and recall of $0.941\pm 0.027$ and $0.930\pm 0.041$ respectively, across all 12 participants.
Figure 4. Normalized confusion matrix of the leave-one-participant-out evaluation across 12 participants of the user study of MunchSonic.
@@ 6. Discussion
6.1. Analyzing the Passive Sensed Ultrasonic Range for Tracking Dietary Actions
@@
Theoretically, our system focuses on nearly inaudible sound frequencies above 18 KHz, which should not interfere with most daily activities, including eating. However, some sounds generated by eating can potentially have frequency components above 18 KHz. To evaluate the impact of these passively sensed ultrasonic signals on distinguishing dietary actions, we conducted a preliminary study with one researcher as a participant. In this study, we turned off both speakers of MunchSonic to prevent them from transmitting any C-FMCW chirps and performed the same set of activities as in the MunchSonic user study. The microphones on both sides continued to record the surrounding acoustic signals while the researcher conducted similar dietary actions. We then created differential echo profile windows (illustrated in the bottom row of Figure 6) in the same way as before and trained the MunchSonic deep learning model. The mean cross-session macro F1-score for this user-dependent model was 0.322. In contrast, when we collected similar data from the same participant with the speakers transmitting ultrasonic chirps, the mean cross-session macro F1-score was 0.976. As illustrated in Figure 6, and evident from the performance difference mentioned, active acoustic sensing plays a crucial role in distinguishing dietary activities by tracking movements of body parts (e.g., face, arms) involved in dietary actions. While passive ultrasonic signals might have some minor impact on detecting fine-grained events, quantifying this impact numerically is challenging and beyond the scope of this paper. Our experiments indicate that chewing crunchy foods (e.g., chips) generated more ultrasonic components in the spectrogram. However, the signal strength of reflections from the ultrasonic chirps transmitted by MunchSonic is much higher than the passively sensed components, potentially overshadowing their feature importance in MunchSonic’s deep learning model training. Further investigation into this will be left for future work.
6.2. Impact of Sensing Range and Temporal Context
@@
We evaluate the impact of the sensing range and temporal context for the MunchSonic system in this experiment. Figure 7 demonstrates that a sliding window size of 2.00 seconds with a 50% overlap yields the best tracking performance. We also observe that a sensing range of 50 to 80 cm demonstrates optimal performance in tracking fine-grained dietary actions. This can be attributed to the fact that most movements related to the dietary actions MunchSonic tracks happen in the jaw and mouth region, which is within 30 cm of the eyeglasses form factor. Additionally, MunchSonic tracks hand movements for food intake, so extending the sensing range to 50 cm yields the best performance. The poorer performance at longer ranges can be rationalized by the fact that various activities can occur in that region, which may not be related to eating.
6.3. Potential Application Scenarios
We envision MunchSonic being potentially useful in various contexts, including eating behavior assessment, chronic disease management, and overall well-being. Our system can enhance eating behavior assessment by providing a passive and objective measurement of eating episodes and micro-level actions (e.g., eating pauses, drinking, chewing) in real-world environments. By accurately classifying these actions, we can derive the ”microstructure of meals” (Bellisle, 2020), including eating duration, eating speed, chewing rate, and chewing efficacy, and their variations across episodes. Traditionally, these metrics are assessed through semi-structured interviews (e.g., Eating Disorder Examination (Fairburn and Wilson, 1993; for Research on Eating Disorders at Oxford, CREDO)) or self-report questionnaires (e.g., (Sasaki et al., 2003)), which rely on recall and approximation and are prone to biases. Our system offers an automated, precise, and naturalistic approach to capturing these behaviors. Clinical studies indicate that these metrics are linked to various health outcomes. For instance, higher eating and chewing rates are associated with higher body mass index (e.g., obesity (Sonoda et al., 2018)), eating disorders (Walsh et al., 1989), and metabolic diseases (Yuan et al., 2021). Therefore, precise assessment of these metrics could enhance the monitoring and management of these health outcomes over time, facilitating real-time monitoring and interventions.
@@ 7. Conclusion
MunchSonic demonstrates accurate tracking of fine-grained dietary actions in real-world settings, surpassing other systems with its single instrumentation point, high precision, privacy preservation, and scalability. Our work holds significant potential for advancing dietary health assessment and monitoring in both technical and healthcare domains.

### v2

@@ 5.1. Evaluation of Frame-Level Inference
We summarize the results of the leave-one-participant-out evaluation of sliding window prediction in Figure 3. According to the evaluation, the average macro F1-score across all 12 participants is 0.935 with a standard deviation of 0.031. It is evident that the MunchSonic system can track fine-grained dietary events with precision and recall of more than 90% for most participants. P08 demonstrates the worst performance in terms of both precision and recall. Additionally, P04 and P13 yielded slightly lower recall compared to the average. Our analysis of the ground truth videos for those participants suggests that the slightly worse performance can be attributed to the participants touching their faces several times during the study for activities not related to eating. This phenomenon is also evident in the confusion matrix in Figure 4. The performance degradation of MunchSonic for face-touch actions can be related to the imbalanced nature of the dataset collected in a completely unconstrained manner, which contains very few samples of this particular activity. Nonetheless, MunchSonic exhibits robust performance across all activities, with a mean precision and recall of $0.941\pm 0.027$ and $0.930\pm 0.041$ respectively, across all 12 participants.
Figure 4. Normalized confusion matrix from the leave-one-participant-out evaluation across 12 participants in the user study of MunchSonic. The values in parentheses represent the total number of instances for each cell.
@@ 6. Discussion
Analyzing the Passive Sensed Ultrasonic Range for Tracking Dietary Actions Theoretically, our system focuses on nearly inaudible sound frequencies above 18 KHz, which should not interfere with most daily activities, including eating. However, some eating sounds can have frequency components above 18 KHz. To evaluate the impact of these passively sensed ultrasonic signals on distinguishing dietary actions, we conducted a preliminary study with one researcher as a participant. In this study, we turned off MunchSonic’s speakers to prevent them from transmitting C-FMCW chirps and performed the same set of activities as in the user study. The microphones continued to record the surrounding acoustic signals. We then created differential echo profile windows and trained the MunchSonic deep learning model. The mean cross-session macro F1-score for this user-dependent model was 0.322. In contrast, with the speakers transmitting ultrasonic chirps, the mean cross-session macro F1-score was 0.976. As illustrated in Figure 6, active acoustic sensing plays a crucial role in distinguishing dietary activities by tracking the movements of body parts involved in dietary actions. While passive ultrasonic signals might have a minor impact on detecting fine-grained events, quantifying this impact numerically is challenging and beyond the scope of this paper. Our experiments indicate that chewing crunchy foods generated more ultrasonic components in the spectrogram. However, the signal strength of reflections from MunchSonic’s transmitted ultrasonic chirps is much higher, potentially overshadowing the passively sensed components’ feature importance. Further investigation into this will be left for future work.
Impact of Sensing Range and Temporal Context We evaluated the impact of sensing range and temporal context for the MunchSonic system. Figure 7 shows that a sliding window size of 2.00 seconds with a 50% overlap yields the best tracking performance. We also found that a sensing range of 50 to 80 cm is optimal for tracking fine-grained dietary actions. This is because most movements related to dietary actions occur in the jaw and mouth region, within 30 cm of the eyeglasses form factor. Additionally, MunchSonic tracks hand movements for food intake, so extending the sensing range to 50 cm yields the best performance. Poorer performance at longer ranges is likely due to various unrelated activities occurring in that region.
@@
Comfort and Safety of Wearing MunchSonic Prototype After participating in the user study described in Section 4, participants completed an IRB-approved questionnaire on the comfort of wearing the MunchSonic prototype. They rated their comfort from 0 (most uncomfortable) to 5 (most comfortable), with a mean rating of 3.462 and a standard deviation of 1.050. Since MunchSonic operates using the ultrasonic range, participants were not expected to hear anything from the system, and indeed, none reported hearing any noise. Three participants noted that the prototype did not fit their head sizes and prescription lenses well, which could be addressed in future iterations by creating an attachable sensing module. Additionally, one participant mentioned that the chest-mounted camera used for ground truth acquisition occasionally touched their dining table, interrupting their eating.
We measured the transmitted signal intensity using a CDC-provided app (Chucri A. Kardous and Peter B. Shaw, [n. d.]), finding it to be 68 dB(A), below the NIOSH limit of 85 dB (Murphy and Franks, 2002). While MHz range ultrasonic exposure can cause discomfort (Moyano et al., 2022), MunchSonic operates in the KHz range with no reported issues. Future studies will explore potential audibility among animals and children.
Potential Application Scenarios We envision MunchSonic being useful for eating behavior assessment, chronic disease management, and overall well-being. Our system passively and objectively measures eating episodes and micro-level actions (e.g., pauses, drinking, chewing) in real-world settings. By accurately classifying these actions, we can derive meal micro-structure metrics (Bellisle, 2020) like eating duration, speed, and chewing rate, traditionally assessed through interviews or self-report questionnaires (Sasaki et al., 2003), which are prone to biases. Our automated approach provides precise, naturalistic data. Clinical studies link these metrics to health outcomes like higher Body Mass Index (Sonoda et al., 2018; Walsh et al., 1989), eating disorders (Fairburn and Wilson, 1993; for Research on Eating Disorders at Oxford, CREDO), and metabolic diseases (Yuan et al., 2021). Accurate assessment enhances monitoring and management of these conditions, facilitating real-time interventions. A day-long evaluation of MunchSonic will provide crucial insights into snacking behavior and bulimia or binge eating disorder detection.
@@ 7. Conclusion
MunchSonic accurately tracks fine-grained dietary actions with high precision and scalability in real-world settings, surpassing other systems and advancing dietary health assessment and intervention design.
Acknowledgements
This project was supported by the National Science Foundation Grant No. 2239569 and partially by the Cornell University IGNITE Innovation Acceleration Program.